OpenStudy (autumnchase):

Give the equation ​(a) in​ slope-intercept form and ​(b) in standard form. m=-3, (7,8)

OpenStudy (natakt):

Any ideas on how you think you would start it?

OpenStudy (natakt):

What would the slope be?

OpenStudy (autumnchase):

m=-3

satellite73 (satellite73):

\[y-y_1=m(x-x_1)\] put \(m=-3,x_1=7,y_1=8\) directly in to the point slope form